OpenStudy (anonymous):

Given the function f(x) = 2(x + 6), find x if f(x) = 22. 56 5 38 17

OpenStudy (jlastino):

22= 2(x+6) Expand 22= 2x+12 22-12=2x 10=2x 5=x